OpenStudy (anonymous):

convert from polar to cartesian: (6, −7π/6)

OpenStudy (jdoe0001):

hint: \(\large \left({\color{blue}{ 6}},{\color{brown}{-\frac{7\pi}{6}}}\right)\qquad \begin{cases} x={\color{blue}{ r}}sin({\color{brown}{ \theta}})\\ y={\color{blue}{ r}}cos({\color{brown}{ \theta}}) \end{cases}\qquad (x,y)\)

OpenStudy (anonymous):

I think you got them mixed up!